Given: S(0)-80, c = S6.00, p-s6.00, K = 78 for both-call and put, r = 5%,at 1 year. Does Put-Call parity holds?

0 A Yes. O B No. Call is undervalued. O C No. Put is overvalued. D No. Either Call is undervalued or Put is overvalued

Your portfolio has 2 options: Call with K-65 and Put with K-70. S(0-$68. Call = $4.70. Put $4.30. Intrinsic value of your portfolio is:
Answer #1

1)

Put call parity equation

C+Ke-rt = S+P

6+78*e-0.05*1 =80+6

80.1 = 86 ---> not holding

D) No, either call is undervalued or put is overvalued

2) value of the portfoilo = value of call + value of put

= (S-K)+(K-S)

= (68-65)+(70-68) = 5